OpenStudy (anonymous):

cylindrical pool with diameter 24 ft and height 4 ft and it was a volume

OpenStudy (callisto):

Volume = base area * height = pi (r)^2 *h = pi (24/2)^2 *4 Can you work it out?

OpenStudy (anonymous):

sure

OpenStudy (callisto):

Glad to hear :)

OpenStudy (anonymous):

is it 576

OpenStudy (callisto):

with a pi follows :D

OpenStudy (anonymous):

sorry not 576 is it was 1808.64. i didn't look at pi

OpenStudy (anonymous):

its 1809.55

OpenStudy (anonymous):

so it was 1809 right

OpenStudy (callisto):

nope round the number to 1810 if you need an integer

OpenStudy (anonymous):

thanks callisto